OpenStudy (anonymous):

write the equation of the line that passes through (4,3) and is parrallel to the line y=x+1

OpenStudy (anonymous):

Since the line has to be parallel to y=x+1, it has to have the same slope, so when you set it up the equation will start like this:\[Y=x+b\]Since y is given as 3 and x equals 4, you know that 3=4+b. Solve for b and you get b=-1

OpenStudy (anonymous):

okay so how would you write it at the end

OpenStudy (anonymous):

as an equation

OpenStudy (anonymous):

y=x-1

OpenStudy (anonymous):

thanks can we do another?